# IXPE Observation of the Low-Synchrotron Peaked Blazar S4 0954+65 During An Optical-X-ray Flare

### v1

@@ 2.1 IXPE observations
IXPE observed S4 0954+65 between 2023 May 25 (MJD 60089) and 2023 June 03 (MJD 60098) with a total exposure time of about 490 ks. We remove the instrumental background following the rejection strategy of Di Marco et al. (2023) and we use the ixpeobssim software (Baldini et al., 2022) to analyse the data. The source events are extracted within a circular region (radius of 1.1 arcmin) centered on the source, while the background events are determined using an annulus with inner and outer radii of 2.0 and 5.0 arcmin, respectively, centered on the source. Figure 1 shows the light curve of the source in the 2–8 keV range obtained using the LC algorithm of ixpeobssim. The source exhibits a flare between MJD 60094 and 60096 with a peak at MJD 60095 (i.e., 2023 May 31).
@@
We determine the polarization properties (in the 2–8 keV energy regime) using the procedure described in Kislat et al. (2015), which is implemented in the PCUBE algorithm of ixpeobssim. Additionally, we obtain the 2–8 keV polarization properties using a spectro-polarimetric fit. The two procedures return consistent results within uncertainties. Henceforth, we focus only on the results of the spectro-polarimetric fit. The energy spectrum is obtained from the PHA1 algorithm of ixpeobsim, while PHA1Q and PHA1U are used to derive spectra for each of the Stokes parameters. Following Di Marco et al. (2022), a weighted analysis is employed to improve the sensitivity. For the total energy spectrum (I-spectrum), the bins are merged with ftgrppha of ftools until a signal-to-noise ratio of at least five is reached. For the spectra of the Stokes parameters (Q- and U-spectra), the binning is chosen to guarantee at least five counts in each bin. The spectro-polarimetric fit is performed via Xspec. To account for the Galactic absorption, we use the Tuebingen-Boulder model (tbabs in Xspec) with an column density of cm-2 (Dickey & Lockman, 1990). We fit the data with a power law model (powerlaw) multiplied by an energy-independent polarization (polconst).
@@
We do not detect X-ray polarization () of S4 0954+65 at a confidence level. When taking the entire IXPE observation window into account, we obtain an upper limit (99.73% confidence level) of . In an attempt to improve the constraints on , we reduce the number of free parameters involved in the spectro-polarimetric fit by assuming a number of physical scenarios: we fix the X-ray polarization angle () to be either parallel or perpendicular to the jet position angle (, see §3.3) or to the measured optical polarization angle (, see §3.3). Additionally, we consider three time intervals: the full IXPE exposure (from MJD 60089 to 60098); the interval in which the optical polarization angle appears to stay constant (MJD 60092 to 60095, see §3); and the period during which the X-ray flux peaks (as detected by IXPE, from MJD 60094 to 60096; see Figure 1). The results of these tests are presented in Table 1. A scenario under which is roughly perpendicular to the constant value of yields the strongest constraints on the X-ray polarization degree, (at 99.73% confidence level). Notably, when is perpendicular to the jet position angle, we obtain (at 99.73% confidence level). We emphasize that fixing other values of typically has a small effect on the upper limits. We generally obtain less constraining upper limits when the shorter time bins are considered.
@@
Since the source exhibits a flare during the IXPE observation, we inspect the data for spectral variability during the observation. All photon spectral index measurements, as calculated with the spectro-polarimetric fit, are consistent with a constant value of . Similarly, all hardness ratio measurements, defined as /) with and being the photon count rates in the 2–4 and 4–8 keV intervals respectively, remain consistent with a constant value of 0.710.01 throughout the IXPE observation, indicating that the source did not any experience significant spectral variations. Therefore, we check for variability in polarization by splitting the IXPE data into both equal and variable time intervals. To obtain the polarization properties in this test, we use both the spectro-polarimetric fit and the PCUBE algorithm. X-ray polarization is not detected in any of the selected time intervals via either of the two methods. Lastly, to check for energy dependence, we divide the IXPE data into two ranges: 2–4 and 4–8 keV. We then apply the spectro-polarimetric fit to both separately, using an energy-independent polarization model (polconst in Xspec). We do not find any significant polarization in either of the two energy bins, as upper limits (at the 3 level) are 20% in all of the tests.
@@ 2.2 XMM-Newton
The XMM-Newton space observatory (Jansen et al., 2001) targeted S4 0954+65 from 2023 May 20 (22:41:00 UT) to May 21 (01:44:53 UT) for a net exposure time of 7.7 ks (Obs. ID 0920901901). The standard routines of the SAS (Science Analysis System, Version 21.0.0) and the latest calibration files are used to calibrate and extract the third-level science products. The background is determined from a circular region with a 40 arcsec radius centered on a source-free area of the CCDs. Then, a circular region with a radius of 22 arcsec is adopted to extract the spectrum of S4 0954+65. This radius is selected based on an iterative process aimed at maximizing the signal-to-noise ratio in the 0.3 to 10 keV regime (e.g., Piconcelli et al. 2004). We subsequently bin the obtained spectrum to have at least 30 counts per bin, while ensuring not to over-sample the instrumental energy resolution by a factor larger than three.
We model the European Photon Imaging Camera pn-CCD (EPIC-pn; Strüder et al. 2001) spectrum with an absorbed, pegged power law (tbabs*pegpwrlw). Additionally, since it is known that there is a strong correlation between the photon index and the absorption parameter, we fix the value of the Galactic hydrogen absorption to (HI4PI Collaboration et al., 2016). We calculate the C-statistics (Cash 1979) of the fit and, in order to assess the goodness of the fit, we additionally compute the theoretical C-statistics as well as its variance, following Kaastra (2017). The best-fit result is given in Table 2.

### v2

@@ 2.1 IXPE observations
IXPE observed S4 0954+65 between May 25, 2023 (MJD 60089) and June 03, 2023 (MJD 60098) with a total exposure time of about 490 ks. We remove the instrumental background following the rejection strategy of Di Marco et al. (2023) and we use the ixpeobssim software (Baldini et al., 2022) to analyse the data. The source events are extracted within a circular region (radius of 1.1 arcmin) centered on the source, while the background events are determined using an annulus with inner and outer radii of 2.0 and 5.0 arcmin, respectively, centered on the source. Figure 1 shows the light curve of the source in the 2–8 keV range obtained using the LC algorithm of ixpeobssim. The source exhibits a flare between MJD 60094 and 60096 with a peak at MJD 60095 (i.e., 2023 May 31).
@@
We determined the polarization properties (in the 2–8 keV energy regime) using the procedure described in Kislat et al. (2015), which is implemented in the PCUBE algorithm of ixpeobssim. Additionally, we obtained the 2–8 keV polarization properties using a spectro-polarimetric fit. The two procedures return consistent results within uncertainties. Henceforth, we focus only on the results of the spectro-polarimetric fit. The energy spectrum was obtained from the PHA1 algorithm of ixpeobsim, while PHA1Q and PHA1U were used to derive spectra for each of the Stokes parameters. Following Di Marco et al. (2022), a weighted analysis was employed to improve the sensitivity. For the total energy spectrum (I-spectrum), the bins were merged with ftgrppha of ftools until a signal-to-noise ratio (S/N) of at least 5 is reached. For the spectra of the Stokes parameters (Q- and U-spectra), a binning was chosen to guarantee at least five counts in each bin. The spectro-polarimetric fit is performed via Xspec. To account for the Galactic absorption, we used the Tuebingen-Boulder model (tbabs in Xspec) with an column density of cm-2 (Dickey & Lockman, 1990). We fit the data with a power law model (powerlaw) multiplied by an energy-independent polarization (polconst).
@@
We did not detect X-ray polarization () of S4 0954+65 at a confidence level. When taking the entire IXPE observation window into account, we obtained an upper limit (99.73% confidence level) of . In an attempt to improve the constraints on , we reduced the number of free parameters involved in the spectro-polarimetric fit by assuming a number of physical scenarios: we fix the X-ray polarization angle () to be either parallel or perpendicular to the jet position angle (, see Sect. 3.3) or to the measured optical polarization angle (, see Sect. 3.3). Additionally, we consider three time intervals: the full IXPE exposure (from MJD 60089 to 60098); the interval in which the optical polarization angle appears to stay constant (MJD 60092 to 60095, see Sect. 3); and the period during which the X-ray flux peaks (as detected by IXPE, from MJD 60094 to 60096; see Fig. 1). The results of these tests are presented in Table 1. A scenario under which is roughly perpendicular to the constant value of yields the strongest constraints on the X-ray polarization degree, (at 99.73% confidence level). Notably, when is perpendicular to the jet position angle, we obtained (at a 99.73% confidence level). We emphasize that fixing other values of typically has a minimal effect on the upper limits. Generally, we obtained less constraining upper limits when the shorter time bins are considered.
@@
Since the source exhibits a flare during the IXPE observation, we inspect the data for spectral variability during the observation. All photon spectral index measurements, as calculated with the spectro-polarimetric fit, are consistent with a constant value of . Similarly, all hardness ratio measurements, defined as /) with and being the photon count rates in the 2–4 and 4–8 keV intervals, respectively, remain consistent with a constant value of 0.710.01 throughout the IXPE observation. This indicates that the source did not experience any significant spectral variations. Therefore, we checked for variability in polarization by splitting the IXPE data into both equal and variable time intervals. To obtain the polarization properties in this test, we used both the spectro-polarimetric fit and the PCUBE algorithm. No X-ray polarization was detected in any of the selected time intervals via either of the two methods. Lastly, to check for energy dependence, we divided the IXPE data into two ranges: 2–4 and 4–8 keV. We then applied the spectro-polarimetric fit to both separately, using an energy-independent polarization model (polconst in Xspec). We did not find any significant polarization in either of the two energy bins, as upper limits (at the 3 level) are 20% in all of the tests.
@@ 2.2 XMM-Newton
The XMM-Newton space observatory (Jansen et al., 2001) targeted S4 0954+65 from 2023 May 20 (22:41:00 UT) to May 21 (01:44:53 UT) for a net exposure time of 7.7 ks (Obs. ID 0920901901). The standard routines of the SAS (Science Analysis System, Version 21.0.0) and the latest calibration files are used to calibrate and extract the third-level science products. The background is determined from a circular region with a 40 arcsec radius centered on a source-free area of the CCDs. Then, a circular region with a radius of 22 arcsec is adopted to extract the spectrum of S4 0954+65. This radius is selected based on an iterative process aimed at maximizing the signal-to-noise ratio in the 0.3 to 10 keV regime (e.g., Piconcelli et al. 2004). We subsequently binned the obtained spectrum to have at least 30 counts per bin, while making sure to avoid oversampling the instrumental energy resolution by a factor larger than 3.
We modeled the European Photon Imaging Camera pn-CCD (EPIC-pn; Strüder et al. 2001) spectrum with an absorbed, pegged power law (tbabs*pegpwrlw). Additionally, since it is known that there is a strong correlation between the photon index and the absorption parameter, we fixed the value of the Galactic hydrogen absorption to (HI4PI Collaboration et al., 2016). We calculated the C-statistics (Cash 1979) of the fit and, to assess the goodness of the fit, we additionally computed the theoretical C-statistics as well as its variance, following Kaastra (2017). The best-fit result is given in Appendix A.